Previous | Calendar |
March |
2016 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
28JD59 |
29JD60 |
1JD61 |
2JD62DB visit (start) fixed sapflow |
3JD63 |
4JD64DB visit (end) |
5JD65 |
6JD66 |
7JD67 |
8JD68 |
9JD69 |
10JD70 |
11JD71Site Visit (SpB): Details , Photos (28) Span cylinder at 800 psi. N2 at 1450 psi from 13:24-13:26 MST, swiped krypton hygrometer at approx 14:00 MST, fixed GPS (wires were pulled out of connector) at 14:20 MST, fixed T9 in NCTC logger (removed TC from Mark cr23x logger) PsT BMW Very warm, sunny and calm. |
12JD72 |
13JD73 |
14JD74 |
15JD75 |
16JD76 |
17JD77snow. |
18JD78more snow. |
19JD79 |
20JD80 |
21JD81 |
22JD82 |
23JD83CU Campus closed A few power outages (in early morning) Heavy snow. |
24JD84 |
25JD85CU Holiday (Spring Break) |
26JD86 |
27JD87 |
28JD88 |
29JD89 |
30JD90Site Visit (HB): Details |
31JD91snow (3-4"). |
1JD92Site Visit (SpB): Details , Photos (27) Span cylinder at 600 psi. N2 at 1000 psi checked on FLIR camera setup SpB Jeep Cool and sunny. |
2JD93 |